to what height must a 2.0 kg box be lifted in order for it to acquire 2.0 joules of energy?
1
Expert's answer
2017-10-13T15:07:07-0400
The potential energy is
E=mgh
The height is
h=E/mg=2.0/(2.0(9.8))=0.10 m.